[Hallc_running] [New Logentry] RC daily update (8/25)

billlee at jlab.org billlee at jlab.org
Tue Aug 25 10:20:08 EDT 2020

Logentry Text:
Dear Collaborators, 

The target (Braina) installation was completed. The hall was locked up ~9pm and beam was established  11pm yesterday. Carbon hole runs was performed. We ran 24uA beam on the reference cell (vacuum) over the night, HMS = -4 GeV at 20 degrees and SHMS = -5.6 GeV at 18 degrees. We collected ~3.5 hours of planned reference (vac) data at this setting and completed it. We will continue with the reference data tonight if the He3 target is not ready for production.

Currently, the accelerator is undergoing a 2 hour repair down time due to the work on 1L13-7 cavity (started at 8:30).  The main objective today is to complete moller measurement starting around 11:00 am. During the moller measurement the target experts will keep their calibration effort for the target. The latest relative target polarization is 37.4% @ 7:42am, note that the real calibration constants are to be determined. 

A couple of cautionary side notes: 1. the beam position on the polarizated target is not yet verified; 2. When rotating the spectrometer, please use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instruction can be found: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT

We will have a 4:15pm meeting. 
Meeting link: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3833987
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200825/04887481/attachment-0001.html>

More information about the Hallc_running mailing list